OpenStudy (anonymous):

if x is inversely proportional to y and x=12 when y=0.4,determine: a.) the value of x when y=3 b.) the value of y when x= 2

OpenStudy (souvik):

x=k/y where k is some constant..put x=12 and y=0.4 and find k...

OpenStudy (anonymous):

:(( huh.. i got confused

OpenStudy (souvik):

12=k/0.4 or k=12*0.4=4.8

OpenStudy (souvik):

now you get x=4.8/y a)put y=3 and calculate x b)put x=2 and calculate y

OpenStudy (anonymous):

thanks!

OpenStudy (souvik):

got your answers?

OpenStudy (anonymous):

therefore the when y=3 then x is1.6 and if x=2 then y is 2.4

OpenStudy (souvik):

well done...

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

 i got the answer of  is it correct???